There are 17 people in your class. What is the probability that two of you share the same birthday?
Answer #1

probability that two of you share the same birthday =1-P(none of 17 share the same birthdays)

=1-(365*364*363*...*(365-17+1))/36517 =1-0.6850 =0.3150